OpenStudy (anonymous):

VERY URGENT. I ONLY NEED ONE THING. What is an example of a non extraneous radical equation!!!!! PLEASE HELP ME

OpenStudy (dangerousjesse):

\[ \sqrt x = \sqrt{-1+x} \sqrt{\frac{x}{-1+x}} \]

OpenStudy (anonymous):

thanks